OpenStudy (anonymous):

Any one knows how to solve sin^-1 y = cos y ? (Not using Computer) Mathematically ?

OpenStudy (anonymous):

1/sin = cos -> 1= sin cos But cos^2 + Sin^2 = 1 Equate, u have a quadratic in Sin or Cos.

OpenStudy (zarkon):

?

OpenStudy (anonymous):

I mean Ssin Inverse !

OpenStudy (zarkon):

you want \[\arcsin(y)=\cos(y)\] I see no way of doing this without some numerical method

OpenStudy (anonymous):

Ah, sorry, thought you meant reciprocal which evidently has no solution (not real ones , anyway).

OpenStudy (zarkon):

sure...as \[1=\sin(\theta)\cos(\theta)\] \[\Rightarrow 2=2\sin(\theta)\cos(\theta)\] \[\Rightarrow 2=\sin(2\theta)\] which is a problem since \[-1\leq\sin(2\theta)\leq1\] :)

OpenStudy (anonymous):

No,I 'm talking about sin Inverse = Cosine

OpenStudy (zarkon):

We know...see my second response.

OpenStudy (anonymous):

No That's wrong

OpenStudy (zarkon):

what is wrong?

OpenStudy (anonymous):

how is sin(x)cos(x)=1 ? For all x ? Just take some values for x as 45 then sin(45)xcos(45)=1/2

OpenStudy (zarkon):

see my 2nd response!..not my 3rd which wasn't directed at you

OpenStudy (zarkon):

i.e. you want \[\arcsin(y)=\cos(y)\] I see no way of doing this without some numerical method

OpenStudy (anonymous):

You see no way of doing it ? Ok !

OpenStudy (zarkon):

it is not possible...Mathematica can't even solve it exactly. It only gives a numerical solution